# Frustrated Random Walks: A Fast Method to Compute Node Distances on Hypergraphs

### v1

@@ III.2 Heavily-weighted and Scale-free hypergraphs
As a second test, we create a hypergraph using the seven Harry Potter novels written by J.K. Rowling. Nodes represent novel characters, and characters that appear in the same scene are collected into the same hyperedge. This hypergraph contains $183$ nodes and $3557$ hyperedges. The maximum node degree is 4548, which is much larger than node number. Thus, this hypergraph is also heavily-weighted. In this hypergraph, the node number is too small for the power law to be obviously visible. One way to see power law in this data set is to expand the hyperedges into edges, and plot the log-log distribution of the edge weights, as shown in Fig. 4.
@@
Figure 4: Log-log plot of edge weight distribution in Harry Potter data set. This figures shows the data set is a heavily-weighted, scale-free hypergraph.
We use SRW, FRW, and DeepWalk to find the nearest neighbors of Harry Potter, and list our results in Table 4. From the table, we can see that the results from FRW and DeepWalk are consistent with human judgment, yet SRW results are not what we expect. This corroborates our assertion that FRW is more suitable than SRW for describing real-world human relationships.
@@
Table 4: Nearest neighbors of Harry Potter according to different methods. The results from FRW and DeepWalk are consistent with human judgement, and beat that from SRW.
IV Time complexity of the expected hitting time method
@@
We can therefore conclude that the time complexity of our algorithm is $O(2E+V)$, with the understanding that $N_{iter}$ is a small constant number. For real-world hypergraphs, we generally have $E\propto V$. In such a case, the time complexity can be simplified to $O(V)$. Thus, the time complexity of our method is approximately linear with respect to the hypergraph size. This linear time complexity for the computation of expected hitting times of random walks is a significant advantage it has compared to the DeepWalk method.
V Comparison of running speeds
@@
VI Conclusion and outlook
In this paper, we introduced the random walk method to calculate node distances on hypergraphs, and benchmark it against DeepWalk. The random walk method includes both simple and frustrated random walks, for both of which the expected hitting times could be exactly computed in a unified framework. We used the expected hitting times of both random walk scenarios to measure node distances on hypergraphs, and tested our methods on several real-world data sets. We showed that for lightly weighted hypergraphs, simple random walk, frustrated random walk and DeepWalk give similar results, whereas for scale-free and heavily weighted hypergraphs, frustrated random walk and DeepWalk are more suitable for calculating node distances. Frustrated random walks can beat DeepWalk in running speeds when the number of target nodes is small. We finally analyzed the time complexity of our random walk method, and showed that it is approximately of linear time complexity.
Although DeepWalk can map hypergraph nodes to vectors, it is an old method that did not take advantage of the emerging graph neural networks. In the future, we want to explore the more efficient hypergraph neural networks to perform node embedding. Hopefully, this can remove our dependence on DeepWalk and accelerate our programs significantly.
@@ Appendix A Computation of expected hitting times on an artificial hypergraph
We use node 3 as the target node, and calculate the expected hitting times of frustrated random walks starting from nodes {0, 1, 2, 4}. First note that, according to our definition in section II.3, node 4 is an adherent to target node 3, since node 4 has the target node as its only neighbor. From Eq. (3), we can see that the transition probability from node 4 to 3 is $p=\frac{1}{2}$. Thus, the expected hitting time from node 4 to 3 is $\mathbb{E}N_{3}^{(4)}=\frac{1}{p}=2$. Next we focus on the calculation of expected hitting times starting from nodes {0, 1, 2}.
@@
To solve Eq. (22), we need the initial condition which is $\boldsymbol{X}_{1}=\begin{pmatrix}0&0&\frac{1}{10}\end{pmatrix}^{T}$. Using the methods outlined in Section II.3.3, we can obtain the probability generating function as
@@
$=\frac{1}{z(16z-65)+50}\begin{pmatrix}z^{2}\\ z^{2}\\ (5-4z)z\end{pmatrix}$ (35)
@@
Table 6: Expected hitting times of frustrated random walks on hypergraph 5 with node 3 as target, starting from nodes {0, 1, 2, 4}, using both analytical and Monte Carlo simulation methods. Analytical methods and Monte Carlo simulation yield consistent results, although analytical method can give high-precision results using much shorter computational time.

### v2

@@ IV.2 Heavily-weighted and scale-free hypergraphs without node labels
We first apply our method to a hypergraph constructed from the seven Harry Potter novels written by J.K. Rowling. This novel was chosen for its intricate plots, complex character relationships and world-wide popularity. In our hypergraph, characters are represented as nodes, and each scene in the novel is represented as a hyperedge. Characters are grouped into the same hyperedge if they appear in the same scene. Node degree indicates the number of times a node (character) appears in the novel. This hypergraph contains $183$ nodes and $3557$ hyperedges. The maximum node degree is 4548, which is much larger than node number. Thus, this hypergraph is heavily-weighted. In this hypergraph, the node number is too small for the power law to be obviously visible. One way to see power law in this data set is to expand the hyperedges into edges, and plot the log-log distribution of the edge weights, as shown in Fig. 4.
@@
Figure 4: Log-log plot of edge weight distribution in Harry Potter data set. This figure shows the data set is a heavily-weighted, scale-free hypergraph.
From our knowledge of the novels, we know that Ron Weasley and Hermione Granger are Harry Potter’s best friends. Thus, a good node-distance-computing algorithm should find these two characters as Harry Potter’s top-2 nearest neighbors, meaning their distances to Harry Potter should be the smallest among all possible nodes. This is the golden criteria we employ in this section to check the validity of our algorithms. We use SRW, FRW, and DeepWalk to rank the nodes by their distances from near to far with respect to Harry Potter, and list our results in Table 4. From the table, we can see that both FRW and DeepWalk list Ron Weasley and Hermione Granger as the nearest neighbors to Harry Potter, yet SRW lists Hokey as the nearest neighbor, which is in sharp contrast with our intuition. This corroborates our assertion that FRW is more suitable than SRW for describing real-world human relationships.
@@
Table 4: Nearest neighbors of Harry Potter (rank order is 1/183) according to different methods. The results from FRW and DeepWalk are consistent with our expectation, meaning Ron Weasley and Hermione Granger are nearest to Harry Potter, and beat that from SRW.
To further compare the results from SRW, FRW and DeepWalk, we list in Table 5 the nearest neighbors of a peripheral character Lucius Malfoy whose rank order by node degree is 46/183. For such a non-central node, the preferential attachment effect which compromises the performance of SRW is much reduced, thus enabling all three methods to yield results that are consistent with human judgement.
SRW FRW DeepWalk
Charity Burbage Narcissa Malfoy Bellatrix Lestrange
Narcissa Malfoy Bellatrix Lestrange Charity Burbage
Andromeda Tonks Augustus Rookwood Augustus Rookwood
Rabastan Lestrange Rodolphus Lestrange Draco Malfoy
Rodolphus Lestrange Rabastan Lestrange Narcissa Malfoy
Augustus Rookwood Theodore Nott Sirius Black
Table 5: Nearest neighbors of Lucius Malfoy (rank order is 46/183) from different methods. All three methods give results consistent with human intuition.
For each target node in the Harry Potter hypergraph, we can use SRW, FRW and DeepWalk to rank the other nodes with respect to the target. To systematically compare the ranking results from SRW and FRW, we use DeepWalk results as our benchmark, and calculate the Spearman correlation coefficients between SRW/FRW ranking results and that from DeepWalk. In Table 6, we show the Spearman correlation coefficients for a list of targets. For the targets to be representative, we select nodes with high, medium and low degrees. From the table, we can see that the ranking results of FRW and DeepWalk are mostly consistent with each other for all the nodes, while SRW can yield results consistent with DeepWalk only for peripheral nodes with low degrees. This is understandable since we have already argued in Section III.2 that in SRW, the minions can have unreasonably small distances with respect to the protagonists, which motivates us to introduce the concept of FRW.
Target Harry Potter Ron Weasley Hermione Granger Voldemort Gilderoy Lockhart Michael Corner Scorpius Malfoy
Node degree 3317 1867 1710 568 77 18 1
SRW -0.0726 0.1266 0.0869 0.2525 0.4957 0.7081 0.5069
FRW 0.5882 0.6517 0.5747 0.5831 0.5485 0.6407 0.4008
Table 6: Results from the Harry Potter data set. Spearman correlation coefficients using DeepWalk results as benchmark, for nodes of high, medium and low degrees. A large coefficient means a high similarity in ranking results with that of DeepWalk.
We then apply our results to a hypergraph created from Dream of the Red Chamber, a novel written by Cao Xueqin. This novel is also famous for its complicated plots and tangled character relationships. Still, we represent characters using nodes and characters appearing in the same scene are grouped into the same hyperedge. This hypergraph contains $324$ nodes and $1,524$ hyperedges. The node degree distribution follows a power law, as shown in Fig 5. The node degrees range from $1$ to $954$ which is much larger than the node number, meaning this hypergraph is also heavily weighted. We will show that for such a hypergraph, the FRW method is most suitable for calculating node distances.
We will again use DeepWalk as our benchmark, and use SRW and FRW to obtain the nearest neighbors for each target node. The Spearman correlation coefficients between SRW/FRW results and DeepWalk results quantify the accuracy of ranking results from SRW and FRW. In Table 7, we show the Spearman correlation coefficients for nodes of various degrees. From this table, we can reach a similar conclusion as that with Harry Potter data set. That is, for central nodes with high degrees, FRW and DeepWalk can yield results that meet our intuition, whereas SRW deviates significantly from either FRW or DeepWalk. Only for peripheral nodes of low degrees can the performance of SRW be on par with that of the other two methods.
Target Jia Baoyu Shi Taijun Wang Xifeng Lin Daiyu Jia Tanchun Jia Yucun Liu Xianglian Wei Ruolan Pan Youan
Node degree 954 697 554 488 188 56 13 1 1
SRW 0.1955 0.246 0.2037 0.4534 0.3193 0.6288 0.4955 0.6913 0.5783
FRW 0.5985 0.577 0.5343 0.6287 0.5823 0.6039 0.463 0.6005 0.5686
Table 7: Spearman correlation coefficients using DeepWalk results as benchmark, for Dream of the Red Chamber data set.
V Time complexity of the expected hitting time method
@@
We can therefore conclude that the time complexity of our algorithm is $O(2E+V)$, with the understanding that $N_{iter}$ is a small constant number. For real-world hypergraphs, we generally have $E\propto V$. In such cases, the time complexity can be simplified to $O(V)$. Thus, the time complexity of our method is approximately linear with respect to the hypergraph size. This linear time complexity for the computation of expected hitting times of random walks is a significant advantage it has over the DeepWalk method.
VI Comparison of running speeds
@@
VII Conclusion
In this paper, we proposed the concept of frustrated random walks on hypergraphs as an improvement of simple random walks and developd a unified framework to compute their expected hitting times. By leveraging the transition matrices of both simple and frustrated random walks, we derived an analytical formula for their expected hitting times and employed the conjugate gradient method or its variants for numerical solutions. We suggested using the expected hitting times of random walks as hypergraph node distances and compared the results with those obtained from DeepWalk.
Our findings indicate that for lightly-weighted hypergraphs, all three methods produce high-quality results. However, for heavily-weighted and scale-free hypergraphs, the simple random walk method yields results significantly deviating from human intuition due to the preferential attachment effect. In contrast, frustrated random walks, which takes into account the preferential attachment effect via the introduction of acceptance probability, provide a more accurate description of such highly complex hypergraphs. Although DeepWalk, originating as a language model, is also effective for computing node distances on these hypergraphs, our method surpasses DeepWalk in terms of running speed by a large margin. Finally, we analyze the time complexity of our random walk method and demonstrate that it is approximately linear for sparse hypergraphs.
@@ Appendix A Computation of expected hitting times on an artificial hypergraph
We use node 3 as the target node, and calculate the expected hitting times of frustrated random walks starting from nodes {0, 1, 2, 4}. First note that, according to our definition in section III.4, node 4 is an adherent to target node 3, since node 4 has the target node as its only neighbor. From Eq. (3), we can see that the transition probability from node 4 to 3 is $p=\frac{1}{2}$. Thus, the expected hitting time from node 4 to 3 is $\mathbb{E}N_{3}^{(4)}=\frac{1}{p}=2$. Next we focus on the calculation of expected hitting times starting from nodes {0, 1, 2}.
@@
To solve Eq. (19), we need the initial condition which is $\boldsymbol{X}_{1}=\begin{pmatrix}0&0&\frac{1}{10}\end{pmatrix}^{T}$. Using the methods outlined in Section III.4.3, we can obtain the probability generating function as
@@
$=\frac{1}{z(16z-65)+50}\begin{pmatrix}z^{2}\\ z^{2}\\ (5-4z)z\end{pmatrix}$
@@
Table 9: Expected hitting times of frustrated random walks on hypergraph 6 with node 3 as target, starting from nodes {0, 1, 2, 4}, using both analytical and Monte Carlo simulation methods. Analytical methods and Monte Carlo simulation yield consistent results, although analytical method can give high-precision results using much shorter computational time.
@@
Appendix B Spectral radius of a substochastic matrix
In this section, we will prove heuristically that the spectral radius of a substochastic matrix $B$ of the shape $n\times n$ must be smaller than one. A rigorous proof of this theorem can be found in Ref. () (https://math.stackexchange.com/users/25136/daniel johnson). By definition, $B$ must satisfy the following conditions:
  1. 1.
    The matrix is non-negative, meaning that $B_{ij}\geq 0,\forall i,j$.
  2. 2.
    $0<\sum_{j=1}^{n}B_{ij}\leq 1,\forall i$, and $\min_{i}\sum_{j=1}^{n}B_{ij}<1,\max_{i}\sum_{j=1}^{n}B_{ij}=1$.
  3. 3.
    $B$ is irreducible.
When applying matrix $B$ to column vector $\mathbbm{1}$, we obtain a new vector
$v^{(1)}=B\mathbbm{1}$ (25)
Since $B$ is substochastic, each element of the new vector $v^{(1)}$ is either 1 or a positive number that is smaller than 1, which we denote as $v^{(1)}<\mathbbm{1}$. Here, for two vectors $\boldsymbol{a},\boldsymbol{b}$ of the same length, $\boldsymbol{a}<\boldsymbol{b}$ means $a_{i}\leq b_{i},\forall i$ and $a_{i}<b_{i}$ exactly holds for at least one $i$. In this section, we will interpret ``<"``"``<"` ` < " (and ``>"``"``>"` ` > ") in this manner whenever it appears between two vectors. We thus claim that matrix $B$ describes a random walk that has a sink. When we apply it to a vector each element of which is positive (we will call it a positive vector), we will lose something from the vector, meaning the vector elements will decrease somehow. Applying $B$ consecutively to $\mathbbm{1}$ yields a series of vectors:
$v^{(k+1)}=Bv^{(k)},\forall k\geq 1$ (26)
From the substochastic nature of $B$, we have
$\mathbbm{1}>v^{(1)}>\ldots>v^{(k)}>v^{(k+1)}>\ldots$ (27)
Since the random walk described by $B$ has a sink that devours everything, we eventually have
$\lim_{k\rightarrow\infty}v^{(k)}=\lim_{k\rightarrow\infty}B^{k}\mathbbm{1}=0$ (28)
That $B$ is non-negative means $B^{k}$ is also non-negative for any integer $k$. Denoting $B^{\infty}=\lim_{k\rightarrow\infty}B^{k}$, we know from $B^{\infty}\mathbbm{1}=0$ that
$B^{\infty}=0$ (29)
This means the spectral radius of $B$ is smaller than 1, just as we expected.
Appendix C Radius of convergence of probability generating function
In this section, we are going to determine the radius of convergence of the probability generating function defined in Eq. (13), which is
$\boldsymbol{f}(z)$ $=\sum_{n=1}^{\infty}\boldsymbol{X}_{n}z^{n}$ (30)
$=\Big{(}\sum_{n=1}^{\infty}B^{n-1}z^{n}\Big{)}\boldsymbol{X}_{1}$
The determination of radius of convergence of function $\boldsymbol{f}(z)$ is tantamount to finding the values of $z$ such that the series $S=\sum_{n=1}^{\infty}B^{n-1}z^{n}$ converges. For frustrated random walks, the coefficient matrix $B$ is real and symmetric, which guarantees that $B$ is diagonalizable and all its eigenvalues are real. In the case of simple random walks, the $B$ matrix is asymmetric and may not be diagonalizable, yet due to the fact that diagonalizable matrices are dense in the set of all matrices, we can still safely assume that $B$ can be diagonalized, although its eigenvalues may be complex. Thus, for an arbitrary matrix $B$, we can assume with probability 1 that there always exists an invertible matrix $U$ and a diagonal matrix $\Lambda$ such that
$B=U\Lambda U^{-1}$ (31)
Here, the diagonal matrix $\Lambda$ has all the eigenvalues of $B$ as its diagonal elements, that is, $\Lambda=diag(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})$, where $N$ is the size of $B$ and $\lambda_{i},i=1,2,\ldots,N$ are its eigenvalues. With the knowledge of this, we can rewrite the series $S$ as
$S$ $=\sum_{n=1}^{\infty}U\Lambda^{n-1}U^{-1}z^{n}$ (32)
$=U\Big{(}\sum_{n=1}^{\infty}\Lambda^{n-1}z^{n}\Big{)}U^{-1}$
$=U\begin{pmatrix}\sum_{n=1}^{\infty}\lambda_{1}^{n-1}z^{n}&&\\ &\ddots&\\ &&\sum_{n=1}^{\infty}\lambda_{N}^{n-1}z^{n}\end{pmatrix}U^{-1}$
For an arbitrary eigenvalue $\lambda_{i}$ which may be a complex number, in order for the series $\sum_{n=1}^{\infty}\lambda_{i}^{n-1}z^{n}$ to converge, we must have $|\lambda_{i}z|<1$, or equivalently,
$|z|<\frac{1}{|\lambda_{i}|}$ (33)
If $\lambda_{i}=0$, we shall interpret the above inequality as $|z|<\infty$.
The inequality (33) should hold for any of the eigenvalues of $B$, which means
$|z|<\frac{1}{\max_{1\leq i\leq N}|\lambda_{i}|}=\frac{1}{\rho(B)}$ (34)
We have thus pinpointed the radius of convergence of the probability generating function at $\frac{1}{\rho(B)}>1$.